# Anticipating Gaming to Incentivize Improvement: Guiding Agents in (Fair) Strategic Classification

### v1

@@ Abstract
As machine learning algorithms increasingly influence critical decision making in different application areas, understanding human strategic behavior in response to these systems becomes vital. We explore individuals’ choice between genuinely improving their qualifications (“improvement”) vs. attempting to deceive the algorithm by manipulating their features (“manipulation”) in response to an algorithmic decision system. We further investigate an algorithm designer’s ability to shape these strategic responses, and its fairness implications. Specifically, we formulate these interactions as a Stackelberg game, where a firm deploys a (fair) classifier, and individuals strategically respond. Our model incorporates both different costs and stochastic efficacy for manipulation and improvement. The analysis reveals different potential classes of agent responses, and characterizes optimal classifiers accordingly. Based on these, we highlight the impact of the firm’s anticipation of strategic behavior, identifying when and why a (fair) strategic policy can not only prevent manipulation, but also incentivize agents to opt for improvement.
footnotetext: Department of Integrated Systems Engineering, The Ohio State University, alhanouti.1@osu.edu. Department of Electrical and Computer Engineering, University of California, San Diego, parinaz@ucsd.edu.
@@ 1 Introduction
Machine learning (ML) algorithms have come to play a pivotal role in guiding decision making in many application areas, including banking, hiring, social media, and resource allocation. While the use of ML-driven systems can enhance efficiency, it can also drive the humans who are subject to algorithmic decisions to adjust their behavior accordingly. Examples include Uber drivers coordinating their behavior in response to its surge pricing algorithm [Möhlmann and Zalmanson, 2017], applicants selecting keywords and formatting to pass automated resume screening [Forbes, 2022], and Facebook users adjusting their posting and content interaction choices in response to the platforms’ curation algorithms [Eslami et al., 2016]. These can be viewed as strategic responses by rational human subjects in these systems, motivating a game-theoretical analysis of learning algorithms with human in the loop.
Earlier works on the study of strategic humans facing ML systems largely focused on scenarios where users can strategically alter only their observable data (e.g., students cheating to obtain better test scores, job applicants making formatting or wording changes to their CV, or loan applicants opening several new accounts to increase their credit scores) to receive a favorable decision (e.g., be accepted to a school, job opening, or loan); see, e.g., [Hu et al., 2019, Milli et al., 2019]. This strategic behavior is referred to as strategic manipulation, where agents change their features without changing their true qualification states. This can be interpreted as cheating the machine learning algorithm: such agents may appear to be more qualified, without being truly suitable for a favorable outcome.
However, strategic responses by human agents need not be limited to manipulation: some agents may choose to put in the effort to truly change their qualification state, instead of simply cheating the system; examples include students studying for (standardized) tests to attain high scores, or loan applicants working on their finances to improve their ability to repay a loan.
This type of strategic behavior, which changes the (observable) features as well as the (unobservable) true qualification state in an attempt to become qualified for a favorable outcome is referred to as strategic improvement; this aspect has been receiving increased attention in the subsequent works on strategic machine learning, e.g., [Haghtalab et al., 2020, Miller et al., 2020, Kleinberg and Raghavan, 2020, Efthymiou et al., 2025]. Generally, from the viewpoint of any institution, the preferred strategic behavior is for users to invest in improvement rather than to try to manipulate their system. Therefore, it is of interest to understand the incentives of strategic agents to opt for each type of response, its implications for the institution’s choice of a decision policy, and potential ways to shape users’ strategic behavior. In this paper, we explore such questions.
In particular, we consider a strategic classification problem, where strategic agents facing a binary classier can decide between exerting effort to either manipulate and only change their features, or improve by changing both their feature and their true qualification state, with the goal of being selected by the classifier. We allow these actions to have different costs, and also differ in the resulting change in the agents’ initial features, with the efficacy of each action being stochastic; as we detail in Section 2, these stochastic efficacies are one of the distinctions of our model from prior works that have accounted for both manipulation and improvement actions in strategic classification. Our goal is to understand the effects of these different costs and efficacies on the agents’ choice to opt for each decision, as well as understand how a firm can leverage this understanding to modify its algorithms to encourage improvements while discouraging manipulations.
To this end, we formulate this setting as a Stackelberg game where the firm first deploys a (fair) classifier, and then agents strategically respond to it. We first characterize the Stackelberg game equilibrium (Proposition 2) under the assumption that improvement is more costly than manipulation, but that it is more effective at advancing agents’ chances of receiving a favorable decision (we do the latter by assuming that the transition probabilities under manipulation first-order stochastically dominate those under improvement). We then show (Proposition 8) how the firm’s knowledge of the strategic behavior (or lack thereof) impacts its choice of a classifier, as well as agents’ strategic choices and outcomes across different qualification states. The main challenge in finding these optimal classifiers is that the regions in which agents opt for the improvement and manipulation actions are themselves functions of the firm’s decision variables, and these dependencies should be accounted for when evaluating the firm’s best-response. This is done by applying the Leibniz integral rule to the firm’s objective function (with adaptations when the firm’s optimal classifier selection problem is subject to fairness constraints), and then leveraging the relation between the threshold and the indifference points in agents’ strategic responses, which appear in the integration limits and their derivatives, to simplify the resulting expressions of the firm’s problem’s optimality conditions.
@@
Building on the characterization of the firm’s optimal policies, we then provide additional insights on the impacts of anticipating agent’s strategic behavior, and the impacts of fairness interventions, on shaping the agents’ responses. Specifically, in Section 6.1, we first discuss how anticipating strategic behavior allows the firm to not only curb undesired manipulation (as also found in prior work) but can also incentivize all agents, whether initially qualified or unqualified, to opt for improvement decisions instead. Notably, we find that the firm will make its selection algorithm more “strict”, incentivizing improvement by unqualified agents (driving them to improve both their qualification states and observable features), while still leaving the manipulation option open to qualified agents (who do not have sufficiently high observable features to be selected otherwise, but whose acceptance benefits the firm). Together, these findings lead to our first key takeaway: anticipating agents’ strategic responses, which the firm does solely with the goal of optimizing its utility, has the added benefit of guiding agents’ behavior in a desirable direction.
Then, in Section 6.2, we investigate the implications of adopting algorithmic fairness interventions, when both improvement and manipulation actions are available to the participating agents. We argue that in its attempt to attain fair outcomes, a firm may decrease improvement incentives for agents from a disadvantaged population; this can be viewed as a negative consequence of fairness interventions on shaping agents’ strategic responses, and emerges even if the differences between demographic groups are only historical (in qualification rates and feature distributions) and not current (in access to strategic resources). As we elaborate through numerical experiments in Section 7, this negative by-product is more severe if the firm is non-strategic (i.e., unaware of agents’ strategic responses). In comparison, a strategic firm recognizes that fairness goals can be met by a combination of adjusting decision thresholds and incentivizing improvement actions. In fact, we observe that fairness desiderata can be achieved with very small changes to the decision thresholds when accounting for agents’ strategic responses, therefore leading to only minor changes in agents’ improvement incentives, and far less loss in utility for the firm. Together, these findings lead to our second key takeaway: existing literature on fairness interventions (e.g., Liu et al. [2018], Hu et al. [2019], Estornell et al. [2023], Keswani and Celis [2023]) often find that fairness is improved in two main ways: inclusiveness (flipping disadvantaged group predictions to positive) or selectivity (flipping advantaged group predictions to negative); we identify a third lever: anticipating strategic responses, which increases improvement incentives, and in turn aids in improving fairness “for free”.
@@
To summarize, our findings offer two new insights for algorithm designers, as well as for managers contemplating the deployment of algorithmic decision rules within their organizations, in the presence of strategic users. First, anticipating the users’ strategic responses to algorithmic systems (which are increasingly observed, as noted in the examples in the beginning of this section) can not only help protect the planned operating points of these systems, but also will ultimately help enhance the system by increasing the number of agents (workers, employees, students, etc.) who select the preferred improvement actions over undesired manipulation actions, in turn leading to an increase in users’ qualification/success levels. Second, as developing fair and responsible algorithmic decision systems is becoming of increased interest due to reputation effects and applicable laws, driving agents’ improvement actions can be viewed as an additional lever in meeting these fairness goals, allowing firms to attain fairness desiderata without having to lower accuracy significantly.
In the remainder of the paper, we begin with a review of related work in Section 2. We then introduce our proposed strategic classification model in Section 3. In Section 4, we characterize agents’ strategic responses under this model, and the resulting post-strategic population statistics, given a fixed classifier. We then characterize the firm’s optimal choice of classifier in Section 5, when the firm is aware or unaware of the impacts of its choice on agents’ responses, as well as with and without fairness interventions. We theoretically analyze the effects of anticipating strategic behavior and the inclusion of fairness constraints in Section 6, followed by numerical experiments and illustrations supporting our analytical findings in Section 7. We conclude with potential directions for future work in Section 8.
@@ 2 Related Work
“Strategic” machine learning, and in particular, strategic classification and regression problems, wherein ML/AI algorithm designer(s) and/or users (rationally) adjust their decisions in response to others’, has become a subject of growing interest. A majority of the research in this area studies settings where humans adjust their behavior with the goal of receiving more favorable algorithmic decisions, and falls within one of three main categories: (i) agents strategically manipulating observable data (features) to secure favorable outcomes without changing their true qualifications [Cohen et al., 2025, Levanon and Rosenfeld, 2022, Sundaram et al., 2023, Cohen et al., 2024, Lechner et al., 2023, Ahmadi et al., 2023, Shao et al., 2023, Levanon and Rosenfeld, 2021, Zhang et al., 2022, Ahmadi et al., 2021, Ebrahimi et al., 2024] and (ii) agents choosing between manipulation and improvement actions, where improvement changes their true qualification [Jin et al., 2023, Wang et al., 2023, Horowitz and Rosenfeld, 2023, Haghtalab et al., 2020, Jin et al., 2022, Miller et al., 2020, Zhang et al., 2020b, Ahmadi et al., 2022, Kleinberg and Raghavan, 2020, Bechavod et al., 2021, Shavit et al., 2020, Harris et al., 2021, Alon et al., 2020, Bechavod et al., 2022, Xie et al., 2024, Efthymiou et al., 2025] . Our work falls within the third category, and differs from the prior literature mainly in our model of both distinct costs as well as different stochastic efficacy for these strategic actions, as well as our exploration of fairness implications of the availability of these actions. We elaborate on these differences below. Before proceeding, we also note that our work (and the closely related literature) touches upon one set of strategic actions; other strategic actions available to firms/agents include strategic interactions between multiple firms [Narang et al., 2022] or agents themselves [Eilat et al., 2022], agents strategically hiding data from the classifier instead of distorting data [Zhang et al., 2020a], or firms adopting randomized classifiers [Braverman and Garg, 2020, Geary and Gouk, 2025].
Among the works that consider manipulation and improvement actions as those available to agents, Shavit et al. [2020] and Bechavod et al. [2021] model different actions available to agents by defining a set of features that may or not affect the true qualification when they change, and propose a conversion matrix to represent the effect of the actions accordingly. Similarly, Horowitz and Rosenfeld [2023],Wang et al. [2023], and Miller et al. [2020] use causal modeling to differentiate actions available to agents that affect the true qualification versus those that only alter predictions. Additionally, Jin et al. [2023], like our work, distinguish between manipulation and improvement actions, but introduce a causal model where an agent’s endowed pre-response attribute causally influences both observable features and the true qualification, with effects captured by a projection matrix.
In contrast, we consider distinct manipulation and improvement actions, with their effects stochastically following different probability distributions. Emelianov et al. [2022], similar to our work, consider the effect of the actions to be probabilistic, but following a normal distribution around the selected effort; in contrast, our proposed model of stochastic efficacy of actions does not specify the boost distributions.
Lastly, our work is also closely related to those that explore the fairness implications of the availability of strategic actions to users. The existing work in this area (e.g., [Milli et al., 2019, Hu et al., 2019, Zhang et al., 2022, Estornell et al., 2023]) study the fairness implications when agents can only manipulate their features. The impacts of the availability of both types of strategic behavior, to our knowledge, remains unexplored. Our work addresses this gap, providing insights that firms can leverage to adapt their algorithms in a way that attains fairness desiderata, in part, by incentivizing improvement decisions.
An early version of this work appeared in [Alhanouti and Naghizadeh, 2024], where we first proposed our model of strategic classification with manipulation and improvement decisions, and provided a preliminary analysis of the impacts of anticipating agents’ strategic behavior. The current paper provides an expanded analysis of anticipating agents’ responses (Proposition 8), an analysis of fairness implications (Corollary 9, and related discussions), and detailed numerical verification and illustrations (Section 7).
@@ 3 Problem Setting and Preliminaries
We consider a Stackelberg game where the firm (algorithm designer) first announces a classifier, following which the agents respond strategically. In this section, we detail the agents’ and the firm’s characteristics, their actions, and their utilities. Our model is similar to, and extends, those of prior work on strategic classification Zhang et al. [2020b, 2022]. The notation is summarized in Table 1.
Table 1: Summary of main notation.
Notation Description
$s\in\{a,b\}$ Demographic group membership.
$n_{s}\in[0,1]$ Fraction of agents in group $s$.
$x\in\mathbb{R}_{\geq 0}$ Agents’ observable feature pre-strategic behavior.
$y\in\{0,1\}$ Agents’ unobservable, true qualification state pre-strategic behavior.
$\alpha_{s}\in[0,1]$ The qualification rate in group $s$ pre-strategic behaviour.
$G^{y}_{s}(x)$ Feature distribution (pdf) of qualification $y$, group $s$, agents, pre-strategic behavior.
$\hat{x},\hat{y}$, $\hat{\alpha}_{s},\hat{G}_{s}^{y}(x)$ Agents’ feature, qualification state, group qualification rate, and feature distribution, post-strategic behavior.
$w\in\{M,I,N\}$ Agents’ strategic actions: manipulation ($M$), improvement ($I$), and doing nothing ($N$).
$C_{w,s}$ Cost of strategic action $w$ to an agent from group $s$.
$\tau^{y}_{w,s}(b)$, $\mathbb{T}^{y}_{w,s}(b)$ The pdf and CDF of the distribution of the boost $b$ from action $w$ for an agent from group $s$ with true qualification $y$.
[…]
$\theta_{s}$ The firm’s decision threshold for group $s$.
$u_{+}/u_{-}$ The firm’s benefit/cost from true/false positives.
$\mathbb{I}^{y}_{s}/\mathbb{M}^{y}_{s}/\mathbb{N}^{y}_{s}$ Set of agents from group $s$ with qualification $y$ who opt for improvement/manipulation/doing nothing.
The agents.
Consider a population of agents with two types of observable attributes: sensitive features that divide the population into two demographic groups $S\in\{a,b\}$ (e.g., gender dividing population into Male and Female, or race dividing the population into White and Non-White agents), and an (observable) feature/score $x\in\mathbb{R}_{\geq 0}$ (e.g., a credit score, an exam score). Let $n_{s}$ be the fraction of agents in group $s$.
Each agent has a true (hidden) label or qualification state, denoted $Y\in\{0,1\}$, with $Y=1$ denoting a qualified agent (e.g., one who can pay off a loan if granted) and $Y=0$ denoting an unqualified agent. Let $\alpha_{s}:=\mathbb{P}(Y=1|S=s)$ denote the qualification rate in group $s$. In addition, let $G^{y}_{s}(x):=\mathbb{P}(X=x|Y=y,S=s)$ denote the probability density function (pdf) of the distribution of the features for individuals with qualification state $y$ from group $s$. We make the following assumptions on the feature distributions.

### v2

@@ Abstract
As machine learning algorithms increasingly influence critical decision making in different application areas, understanding human strategic behavior in response to these systems becomes vital. We explore individuals’ choice between genuinely improving their qualifications (“improvement”) vs. attempting to deceive the algorithm by manipulating their features (“manipulation”) in response to an algorithmic decision system. We further investigate an algorithm designer’s ability to shape these strategic responses, and its fairness implications. Specifically, we formulate these interactions as a Stackelberg game, where a firm deploys a (fair) classifier, and individuals strategically respond. Our model incorporates both different costs and stochastic efficacy for manipulation and improvement. The analysis reveals different potential classes of agent responses, and characterizes optimal classifiers accordingly. Based on these, we identify when and why a (fair) strategic policy can not only prevent manipulation, but also incentivize agents to opt for improvement. Our findings shed light on the intertwined nature of microeconomic and ethical implications of firms’ anticipation of strategic behavior when employing ML-driven decision systems.
footnotetext: 1Department of Integrated Systems Engineering, The Ohio State University, alhanouti.1@osu.edu. 2Department of Electrical and Computer Engineering, University of California, San Diego, parinaz@ucsd.edu. 3Department of Industrial Engineering, Jordan University of Science & Technology.
@@ 1 Introduction
Machine learning (ML) algorithms have come to play a pivotal role in guiding decision making in many application areas, including banking, hiring, social media, and resource allocation. While the use of ML-driven systems can enhance efficiency, it can also drive the humans who are subject to algorithmic decisions to adjust their behavior accordingly. For instance, candidates in school admissions, job recruitment, and research pools tend to systematically shift their self-presentation to appear more analytical when facing algorithmic assessments [Goergen et al., 2025]. In other algorithmic contexts, Uber drivers coordinate their behavior to trigger surge pricing [Andon, 2023, Möhlmann and Zalmanson, 2017], and users change their posting and engagement behavior to influence content curation [Cen, 2024, Eslami et al., 2016]. Such strategic responses motivate a game-theoretic analysis of learning algorithms with humans in the loop.
Existing work on strategic classification have primarily focused on agents who manipulate observable features to obtain a favorable prediction without changing their true qualification (e.g., opening credit lines, cosmetically editing resumes, or cheating to obtain better test scores). More recent works (e.g., [Efthymiou et al., 2025, Xie et al., 2024b, Tsirtsis et al., 2024, Wang et al., 2023, Horowitz and Rosenfeld, 2023]) have further incorporated the possibility of improvement: individuals can invest effort to genuinely improve their qualification state, by, e.g., studying for exams or building financial stability. From the perspective of a decision-making institution, improvement is strictly preferable to manipulation, raising the question of when and how a classifier can incentivize improvement rather than gaming by its strategic users.
This paper examines such incentives in a strategic classification setting in which agents may choose between manipulation and improvement. We allow these actions to differ both in cost and in the stochastic efficacy of feature changes; the later is an aspect that, unlike prior deterministic treatments of strategic responses, plays a central role in our analysis (see Section 2). We formulate the interaction as a Stackelberg game: the firm selects a (possibly fairness-constrained) classifier, and agents best-respond strategically. In Proposition 2, we show that, under mild assumptions (improvement is costlier but more beneficial in expectation), equilibrium behavior depends on how the firm’s threshold interacts with agents’ cost and boost distributions. Proposition 8 then uses these findings to characterize how a strategic firm adjusts its policy relative to a non-strategic one. The main technical challenge is that the regions in which agents choose manipulation or improvement depend endogenously on the firm’s decision variable, requiring the firm’s objective to be evaluated over moving integration limits; we address this using the Leibniz integral rule (with modifications for fairness constraints) and the linkage between thresholds and the agents’ indifference points to obtain tractable optimality conditions.
@@
Building on this characterization, we find that a strategic firm not only curbs manipulation (as also found in prior work) but also adopts stricter selection thresholds that encourage all agents, whether initially qualified or unqualified, to opt for improvement decisions instead. Specifically, incentivizing improvement by unqualified agents (driving them to improve both their qualification states and observable features), while still leaving the manipulation option open to qualified agents (who do not have sufficiently high observable features to be selected otherwise, but whose acceptance benefits the firm). Together, these findings lead to our first key takeaway: anticipating agents’ strategic responses, which the firm does solely with the goal of optimizing its utility, has the added benefit of guiding agents’ behavior in a desirable direction.
Then, in Section 6.1, we study the effects of fairness interventions when both manipulation and improvement are available. We argue that, in attempting to attain fair outcomes, a firm may inadvertently decrease improvement incentives for disadvantaged groups. This negative by-product arises even when group disparities are purely historical (in qualification rates or feature distributions) rather than current differences in access to strategic resources. Our numerical results in Section 7 further show that fairness interventions, when imposed by a non-strategic firm, can unintentionally reduce improvement incentives for disadvantaged groups despite equalizing decision rates. In contrast, a strategic firm can satisfy the same fairness criteria while preserving (and even amplifying) improvement incentives, by jointly adjusting thresholds and leveraging endogenous behavioral responses. These findings point out a third lever for fairness beyond inclusiveness (flipping disadvantaged group predictions to positive) and selectivity (flipping advantaged group predictions to negative) [Estornell et al., 2023, Keswani and Celis, 2023, Hu et al., 2019, Liu et al., 2018, Zafar et al., 2017]: anticipating strategic responses to induce improvement and in turn improving fairness “for free”. Our findings are supported (and expanded) by numerical experiments and illustrations in Section 7, conducted on semi-synthetic data constructed from real-world FICO credit score statistics.
@@
To summarize, our findings offer two new insights for algorithm designers. First, anticipating the users’ strategic responses to algorithmic systems (which are increasingly observed, as noted in the examples in the beginning of this section) can not only help protect the planned operating points of these systems, but also will ultimately help enhance the system by increasing the number of agents (workers, employees, students, etc.) who select the preferred improvement actions over undesired manipulation actions, in turn leading to an increase in users’ qualification/success levels. Second, as developing fair and responsible algorithmic decision systems is becoming of increased interest due to reputation effects and applicable laws, driving agents’ improvement actions can be viewed as an additional lever in meeting these fairness goals, allowing firms to attain fairness desiderata without having to lower accuracy significantly.
@@ 2 Related Work
Strategic machine learning has received growing attention in settings where agents alter their behavior in order to obtain more favorable algorithmic outcomes. Most existing work falls into two main categories: (i) models in which agents manipulate observable features without changing their underlying qualification [e.g., Cohen et al., 2025, Ebrahimi et al., 2025, Cohen et al., 2024, Ahmadi et al., 2024, Sundaram et al., 2023, Lechner et al., 2023, Ahmadi et al., 2023, Shao et al., 2023, Zhang et al., 2022, Levanon and Rosenfeld, 2021, Ahmadi et al., 2021], and (ii) models where agents may also improve their true qualification state through costly investment [e.g., Efthymiou et al., 2025, Xie et al., 2024a, Jin et al., 2023, Wang et al., 2023, Horowitz and Rosenfeld, 2023, Ahmadi et al., 2022, Jin et al., 2022, Haghtalab et al., 2020, Bechavod et al., 2022, Zhang et al., 2020b, Kleinberg and Raghavan, 2020, Bechavod et al., 2021, Harris et al., 2021, Shavit et al., 2020, Alon et al., 2020]. A growing literature also considers fairness implications in these settings, though primarily under manipulation-only behavior [Milli et al., 2019, Hu et al., 2019, Zhang et al., 2022, Estornell et al., 2023]. Our work fits within this second category, but departs from prior approaches in two key respects. First, unlike causal or feature-based formulations that distinguish manipulation from improvement through structural relations among features [Chen et al., 2023, Horowitz and Rosenfeld, 2023, Wang et al., 2023, Shavit et al., 2020, Bechavod et al., 2021, Miller et al., 2020], we explicitly model heterogeneous costs and stochastic efficacy for the two action types. While Emelianov et al. [2022] also introduce stochastic action effects via normally distributed boosts, we impose no parametric restriction on the boost distribution. This modeling choice enables us to analyze not only whether agents choose to improve, but when improvement is actually incentivized by the economic environment.
Our work is also closely related to [Xie et al., 2024b], which model manipulation and improvement within a Stackelberg framework, but assume that only unqualified agents behave strategically by imitating qualified agents. In contrast, we allow all agents to behave strategically without imitation. (We also differ in some other model elements.) Notably, we show that the effects of anticipating strategic behavior differs across qualified and unqualified individuals. In addition, Xie et al. [2024b] also study how anticipating strategic responses affects fairness across demographic groups and, like us, find that fairness coincides with incentivizing improvement and discouraging manipulation. While we consider similar environments and reach aligned high-level conclusions, we differ in how fairness is modeled. Their analysis evaluates fairness through outcome disparities under a given (potentially unfair) policy, whereas we explicitly model fair decision-making as a constrained optimization problem, deriving policies that satisfy fairness constraints by construction and enable analyses of different fairness interventions that are not accessible under outcome-based fairness evaluation alone.
Finally, we note that there are other forms and dimensions of strategic responses considered in the literature, but that are outside our scope; these include strategic coordination across agents or firms [Narang et al., 2022], strategic hiding [Zhang et al., 2020a], behavioral biases [Ebrahimi et al., 2025], and randomized policies [Geary and Gouk, 2025, Sundaram et al., 2023, Braverman and Garg, 2020].
An early version of this work appeared in [Alhanouti and Naghizadeh, 2024]; the present paper provides a full characterization of the firm’s optimal response (Proposition 8), the fairness–incentive tradeoffs it induces (Corollary 9), and extensive empirical validation on semi-synthetic dataset based on real-word FICO credit data (Section 7).
@@ 3 Problem Setting and Preliminaries
We consider a Stackelberg game where the firm (algorithm designer) first announces a classifier, following which the agents respond strategically. In this section, we detail the agents’ and the firm’s characteristics, their actions, and their utilities. Our model is similar to, and extends, those of prior work on strategic classification (e.g., Zhang et al. [2020b, 2022]).
3.1 The agents
Consider a population of agents with two types of observable attributes: a feature/score $x\in\mathbb{R}_{\geq 0}$ (e.g., a credit score, an exam score) and a sensitive feature that divide the population into two demographic groups $S\in\{a,b\}$ (e.g., sex dividing population into Male and Female, or race dividing the population into White and Non-White agents). Let $n_{s}$ be the fraction of agents in group $s$.
Each agent has a true (hidden) label or qualification state, denoted $Y\in\{0,1\}$, with $Y=1$ denoting a qualified agent (e.g., one who can pay off a loan if granted) and $Y=0$ denoting an unqualified agent. Let $\alpha_{s}:=\mathbb{P}(Y=1|S=s)$ denote the qualification rate in group $s$. In addition, let $G^{y}_{s}(x):=\mathbb{P}(X=x|Y=y,S=s)$ denote the probability density function (pdf) of the distribution of the features for individuals with qualification state $y$ from group $s$. We make the following (common) assumption on the feature distributions.